OpenStudy (anonymous):

Line JK contains points J (4, 3) and K (1, 5). Line LM contains points L (2, 3) and M (−1, 5). Lines JK and LM are -Parallel -Perpendicular -Neither

OpenStudy (anonymous):

@perl

OpenStudy (perl):

lines are parallel if they have the same slope

OpenStudy (anonymous):

For the slopes I got 2/-3 and 2/3. Did I do it right?

OpenStudy (perl):

Line JK contains points J (4, 3) and K (1, 5). Line LM contains points L (2, 3) and M (-1, 5). slope JK = (5-3)/(1-4) slope LM= (5- 3)/(-1-2)

OpenStudy (anonymous):

ohhh so they are parallel! Thank you!!